### v2

# Decoding Open-Ended Information Seeking Goals from Eye Movements in Reading

Abstract
When reading, we often have specific information that interests us in a text. For example, you might be reading this paper because you are curious about LLMs for eye movements in reading, the experimental design, or perhaps you wonder “This sounds like science fiction. Does it actually work?”. More broadly, in daily life, people approach texts with any number of text-specific goals that guide their reading behavior. In this work, we ask, for the first time, whether open-ended reading goals can be automatically decoded solely from eye movements in reading. To address this question, we introduce goal decoding tasks and evaluation frameworks using large-scale eye tracking for reading data in English with hundreds of text-specific information seeking tasks. We develop and compare several discriminative and generative multimodal text and eye movements LLMs for these tasks. Our experiments show considerable success on the task of selecting the correct goal among several options, and even progress towards free-form textual reconstruction of the precise goal formulation. These results open the door for further scientific investigation of goal driven reading, as well as the development of educational and assistive technologies that will rely on real-time decoding of reader goals from their eye movements. [^1] Code is available anonymously here.
1 Introduction
Eye movements in reading are a key methodology for studying the cognitive basis of reading and language processing. When we read, our eyes move over the text in a saccadic manner, with prolonged periods of time in which the gaze is stable, called fixations, and fast transitions between fixations called saccades (Schotter & Dillon, 2025). This trajectory contains rich behavioral traces of the ways in which readers interact with texts and process language in real time (Rayner, 1998; Just & Carpenter, 1980). Understanding the nature and strength of the relation between eye movements, the text, and online language processing has been a major research avenue in the psychology of reading in the past few decades, and in recent years has also been gaining increasing interest in NLP and ML (Barrett & Hollenstein, 2020; Reich et al., 2025).
In this work, we investigate a reading scenario that is prevalent in daily life but remains understudied – seeking specific information in a text. This scenario deviates from a common implicit assumption in psycholinguistics, according to which the comprehender’s goal is constant across communicative contexts: a general understanding of the linguistic input. In practice, readers approach texts with a variety of goals and information seeking needs. Each such goal can have a profound impact on the cognitive processes of language comprehension and the corresponding behavioral traces of eye movements over the text (Radach & Kennedy, 2004; Kaakinen & Hyönä, 2010; Hahn & Keller, 2023; Shubi & Berzak, 2023).
Our study focuses on the following question: can arbitrary, text-specific information goals be automatically decoded from eye movements in reading? Specifically, given a single participant reading a single passage with a question in mind that they would like to answer via the text, can this question be decoded from their eye movements over the passage? This task has not been addressed in prior research and has both scientific and practical importance. Scientifically, it allows obtaining insights into the strength and nature of the relation between task conditioning and reading behavior, and more broadly, the extent to which eye movements contain information on the cognitive state of the reader. Practically, it opens the door for applications in education, content personalization, and text accessibility, which will rely on detecting information seeking behavior, its specific purpose, and the extent to which this behavior is effective.
Our key contributions are the following:
Tasks: We present a new cognitive state decoding challenge from eye movements in reading: decoding arbitrary information seeking goals over the content of specific texts. We instantiate this challenge as a goal selection task given several possible goals, and as an even more challenging open-ended goal reconstruction task where the information-seeking goal has to be generated.
Models: We introduce two strong baseline models and adjust two state-of-the-art discriminative models that combine text with eye movements for the goal selection task. We further develop two new generative text and eye movements multimodal LLMs that can be used for both goal selection and goal reconstruction.
Evaluations: We perform systematic evaluations of selection accuracy against informative baseline models at two levels of task difficulty and utilize several generation quality measures for the reconstruction task. For both tasks, we examine model generalization across new texts, new readers, and the combination of both.
2 Eye Tracking Data
We use OneStop (Berzak et al., 2025), a public eye tracking dataset collected with a high-end EyeLink 1000 Plus eye tracker. It comprises 360 adult native English speakers reading Guardian articles from the OneStopQA multiple choice reading comprehension dataset (Berzak et al., 2020; Vajjala & Lučić, 2018). In each experimental trial, a participant reads a paragraph on a screen and, on the following screen, answers a multiple-choice reading comprehension question, without the ability to return to the paragraph. 180 participants read in an information seeking condition, where they receive the question (but not the answers) before reading the paragraph, and therefore have a specific reading goal for the upcoming paragraph. The remaining 180 participants do not receive the question ahead of time, and thus have to be ready to answer any subsequent question. Here, we use the information seeking part of OneStop, which is uniquely suited for our study.
Figure 1: Example of a text from OneStop, consisting of a paragraph and three questions. Two of the questions have the same critical span: the paragraph segment essential for answering the question.
For each paragraph, a participant receives one of three possible questions over the paragraph. Each such question has a manually annotated critical span in the paragraph, which contains the information that is essential for answering the question correctly. Two of the questions share the same critical span, while the third question has a distinct critical span. This design supports a question selection task with two tiers of difficulty, distinguishing between questions with different critical spans and the more challenging variant of distinguishing between questions that share the same critical span. Figure˜1 presents an example of a paragraph, its three questions, and their corresponding critical span annotations.
The data has 486 questions over 162 paragraphs, with 20 participants answering each question. The mean question length is 10 words. The mean paragraph length is 109 words (5.8 sentences), of which the critical span is 32 words. Overall, there are 1,055,429 paragraph word tokens for which eye tracking data was collected in the information seeking regime.
3 Problem Formulation
The general problem we address is using eye movements over a text (in our setting, a text is always a paragraph) to predict which question a participant received prior to reading that text. Each text $T$ has three text specific questions, $Q_{T}=\left\{q_{1,c_{1}},q_{2,c_{2}},q_{3,c_{2}}\right\}$. The critical information for answering $q_{1,c_{1}}$ is located in the text in span $c_{1}$, while for both $q_{2,c_{2}}$ and $q_{3,c_{2}}$ it is in span $c_{2}$. In each experiment trial, a participant $P$ is presented with one question, $q_{P}\in Q_{T}$, before reading the text. The recording of the participant’s eye movements while reading the text is denoted by $E_{P,T}$.
We propose two task variants as follows.
Question Selection In the selection task, a classifier $h$ is trained to select which question from $Q_{T}$ the participant received, from their eye movement recording over the text:
$h(T,Q_{T},E_{P,T})\longrightarrow\hat{q}_{P}\in Q_{T}$
Question Reconstruction In the free-form reconstruction task, a generative model $g$ is trained to generate the question that the participant received from their eye movement recording over the text:
$g(T,E_{P,T})\longrightarrow\hat{q}_{P}$
4 Models
We introduce discriminative and generative models, presented schematically in Figure˜2.
(a) Discriminative models (Section˜4.1)
(b) Generative models (Section˜4.2)
Figure 2: Two model types for decoding the question presented to a participant before reading the text, from their eye movements while reading the text. (a) Discriminative models that score candidate questions and are evaluated on question selection accuracy. (b) Generative models trained to reconstruct the question presented to the reader, and evaluated on question selection accuracy (via log-likelihood assignment to each candidate question) and free-form question reconstruction quality.
4.1 Discriminative Models
Reading-Time Informed Embedding Similarity Models
We develop two simple baseline models which build on the following observations from prior literature: (i) readers tend to spend more time on question-relevant than on question-irrelevant portions of the text (Hahn & Keller, 2023; Malmaud et al., 2020; Shubi & Berzak, 2023), and (ii) question-relevant text segments tend to have higher semantic similarity to the question compared to question-irrelevant segments (Mitra et al., 2016).
(a) Question Similarity to RT-Weighted Passage
(b) RT Similarity to Question-Word Similarities
Figure 3: Reading-Time Informed Embedding Similarity models for the question selection task.
Question Similarity to RT-Weighted Passage (Figure˜3(a)) For a participant reading a text of $N$ words, we compute a text embedding, $\text{emb}(T)$, as a weighted average of its contextualized word embeddings, where the weights are speed-normalized word reading times, $RT(w_{i})$. Formally: $\text{emb}(T)=\sum_{i=1}^{N}\text{RT}(w_{i})\,\text{emb}(w_{i})$. The embeddings for word $w_{i}$, $\text{emb}(w_{i})$, are extracted from RoBERTa (Liu et al., 2019), and $\text{RT}(w_{i})$ is the word’s Total Fixation Duration (i.e., sum of all the fixation durations) normalized by the text’s total reading time. For each candidate question $q_{j}$, we use its [CLS] token embedding as $\text{emb}(q_{j})$, and compute its cosine similarity with $\text{emb}(T)$. We then predict the question with the highest similarity:
${\arg\max}_{j\in\{1,_{\cdots},|Q_{T}|\}}\cos(\text{emb}(T),\text{emb}(q_{j}))$
RT Similarity to Question-Word Similarities (Figure˜3(b)) An alternative to the model above. Compares the vector of word reading times for the text to the vector of word–question similarities. Specifically, for each candidate question $q_{j}$ and text word $w_{i}$, we first compute $\text{sim}(q_{j},w_{i})=\cos\bigl(\text{emb}(q_{j}),\text{emb}(w_{i})\bigr)$, where as before, $\text{emb}(w_{i})$ is a contextual word embedding from RoBERTa, and $\text{emb}(q_{j})$ is the question [CLS] embedding, yielding a length-$N$ vector of similarities $\text{sim}(q_{j},T)=[\text{sim}(q_{j},w_{1}),\text{sim}(q_{j},w_{2}),...,\text{sim}(q_{j},w_{n})]$. We select the question maximizing the cosine similarity between this vector and $\text{RT}(T)$, the length-$N$ vector of speed-normalized Total Fixation Durations for the text words:
${\arg\max}_{j\in\{1,_{\cdots},|Q_{T}|\}}\cos\bigl(\text{sim}(q_{j},T),\,\text{RT}(T)\bigr)$
State-of-the-art Neural Models
We further adapt two state-of-the-art encoder models for eye movements in reading to our task. Building on the approach used in Radford et al. (2019) for encoding multiple-choice reading comprehension items, we use several copies of each text, where instead of combining each copy with a candidate answer, we combine it with a candidate question, and further provide the eye movements over the text $\langle T,q\in Q_{T},E_{P,T}\rangle$. The models assign a probability to each triplet and select the highest probability question.
Haller RNN (Haller et al., 2022) An RNN-based model that orders word embeddings based on the fixation sequence and encodes each fixation using eye movement features. We adapt this model by appending question embeddings to the input and processing them through the RNN, followed by a classifier scoring each candidate question. Additional details on the model, including a diagram of the modified architecture and the used eye movement features, are presented in Section˜A.1.
RoBERTEye-Fixations (Shubi et al., 2024) A transformer-based model which integrates fixation-level eye movement features into RoBERTa (Liu et al., 2019). An eye movements feature vector for each fixation is first projected to the dimension of the word embeddings. These fixation vectors are then aligned with their corresponding words via the addition of position embeddings and finally concatenated with the word sequence. We adapt this model to incorporate candidate question embeddings and modify the classification layer to score questions. Section˜A.2 provides further details, including a model diagram and the eye movement features that the model uses.
4.2 Generative Models
We introduce two LLM-based approaches for generating the question based on the eye movements over the text. Both models are fine-tuned using an autoregressive next-token prediction objective for the correct question with a cross-entropy loss. During fine-tuning, the language models’ parameters remain frozen, and only an additional LoRA (Hu et al., 2022) adapter component is trained. As described in Section˜5, we evaluate these models both on the question selection and the question reconstruction tasks.
DalEye-LLaVA is based on LLaVA-OneVision (Li et al., 2024), a vision-language model that uses the Qwen 2-0.5B language model (Yang et al., 2024) as a backbone. The model input consists of a task prompt, the text, and eye movement feature embeddings for the text words. We replace the vision encoder of LLaVA-OneVision with a novel trainable eye movement encoder, which encodes eye movement features using fully connected and convolutional layers. The eye movements embeddings are positionally aligned with the text words. Additional details, including a model diagram and eye movement features, are provided in Section˜B.1.
DalEye-Llama In this approach, we provide the task and the eye movement trajectory of a participant over the text to the Llama 3.1 language model (Grattafiori et al., 2024) in textual form. The model is finetuned to reconstruct the true question given this input. We use a fixation-based eye movements representation, where each fixation is represented as a tuple containing the index of the fixated word, the word itself, the fixation duration in milliseconds, and the direction of the next saccade. This encoding preserves both the spatial locations and temporal order of the fixations on the text. It follows a prompt describing the task and instructing the model to generate the question that was presented to the reader. In Section˜B.2, we provide an example of the model input for a complete fixation sequence over a paragraph. We further report experimental results for a differently worded task prompt, as well as two alternative representations of the eye movement sequence, one based on word-level features and another combining word- and fixation-level features.
5 Experimental Setup and Evaluation
5.1 Data Splits
The OneStop eye tracking data is not i.i.d., each participant reads multiple paragraphs, and each paragraph is read by multiple participants. Following prior work on predictive modeling with OneStop (e.g. Shubi et al., 2024), we use a 10-fold cross-validation procedure where each of the 10 data splits has a training, validation, and a test set. To account for the structured nature of the data, the validation set and the test set are each partitioned into three disjoint parts that capture three different levels of model generalization, ordered by task difficulty:
New Participant: Training eye tracking data is available for the text but not for the participant.
New Text: Training eye tracking data is available for the participant but not for the text.
New Text & Participant: Neither the participant nor the text was in the training data.
Each data split includes 64% of the trials in the training set, 17% in the validation set, and 19% in the test set – divided into 9% New Participant, 9% New Text, and 1% New Text & Participant. When aggregated across the 10 splits, 90% of the trials appear in both New Participant and New Text evaluation regimes, and 10% appear in the New Text & Participant regime. We ensure that in each split into training, validation, and test sets, there is a balanced distribution of question types, i.e. $q_{1,c_{1}},q_{2,c_{2}},q_{3,c_{2}}$ for each text in each part of the split. Hyperparameter optimization and model selection are conducted separately for each split. We assume that the evaluation regime with respect to novelty of items and participants at test time is unknown, and therefore model selection is performed using the entire validation set within each split. Further details on the hyperparameter search space, training, and evaluation procedures are provided in Appendix˜C.
5.2 Question Selection
We evaluate all the models on the question selection task and report selection accuracy. As described in Section˜4, the discriminative models assign a probability to a triplet $\langle T,q\in Q_{T},E_{P,T}\rangle$, and select the candidate question with the highest probability. For the generative models, we select the question with the highest model-assigned log-likelihood. In addition to an All overall three-way selection accuracy, we also report a breakdown of the model predictions into two fine-grained evaluations:
Different critical spans This evaluation tests the model’s ability to distinguish between questions with different critical spans, reducing the selection problem to a binary decision between $q_{1,c_{1}}$ and $\{q_{2,c_{2}},q_{3,c_{2}}\}$. Note that random choice accuracy in this evaluation is approximately 55%, rather than 50%, due to a Monty Hall-type setup arising from grouping two questions under one critical span (see Table˜1).
Same critical span This is a more challenging evaluation for differentiating between the two questions for $c_{2}$: $q_{2,c_{2}}$ versus $q_{3,c_{2}}$. Note that here we disregard the question whose critical span is $c_{1}$. The corresponding chance accuracy is 50%.
Table 1: Overview of (a) the three-way question selection task and the breakdown of this task into (b) questions with different critical spans and (c) questions with the same critical span. Chance accuracy under the three evaluation regimes: (a) ${\color[rgb]{0.265625,0.59375,0.1640625}3}/9$ (33.3%), (b) ${\color[rgb]{0.265625,0.59375,0.1640625}5}/9$ (55.5%), (c) ${\color[rgb]{0.265625,0.59375,0.1640625}2}/4$ (50%).
(a) All (b) Different critical spans (c) Same critical span
Predicted $\color[rgb]{0.9765625,0.56640625,0.40625}q_{1,c_{1}}$} $\color[rgb]{0.08203125,0.49609375,0.49609375}q_{2,c_{2}}$} $\color[rgb]{0.08203125,0.49609375,0.49609375}q_{3,c_{2}}$} $\color[rgb]{0.9765625,0.56640625,0.40625}q_{1,c_{1}}$} $\color[rgb]{0.08203125,0.49609375,0.49609375}q_{2,c_{2}}$} $\color[rgb]{0.08203125,0.49609375,0.49609375}q_{3,c_{2}}$} $\color[rgb]{0.9765625,0.56640625,0.40625}q_{1,c_{1}}$} $\color[rgb]{0.08203125,0.49609375,0.49609375}q_{2,c_{2}}$} $\color[rgb]{0.08203125,0.49609375,0.49609375}q_{3,c_{2}}$}
True
$\color[rgb]{0.9765625,0.56640625,0.40625}q_{1,c_{1}}$}
$\color[rgb]{0.08203125,0.49609375,0.49609375}q_{2,c_{2}}$}
$\color[rgb]{0.08203125,0.49609375,0.49609375}q_{3,c_{2}}$}
5.3 Question Reconstruction
The generative models are further evaluated for the quality of the reconstructed questions. As with other generation tasks, evaluating reconstruction quality is a major challenge. Here, we use several heuristic evaluations for both the surface form and the semantic content of the generated questions.
The question word We check if the generated question word matches the question word of the ground truth question. The possible question words are What, When, Where, Who, Whom, Which, Whose, Why, and How. An additional category is Other, reserved for all other words starting a question (7% of the questions). The agreement with the ground truth questions is measured using pairwise Cohen’s Kappa (Cohen, 1960).
UIUC question category We test if the generated question has the same semantic category as the ground truth question, using the question-type taxonomy of Li & Roth (2002). We use GPT-4o (Hurst et al., 2024) to categorize the questions into the main taxonomy categories: entities, humans, numeric, and locations, and the subcategories of ‘descriptions’: manner, reasoning, definition, and description due to their high frequency in the data. We measure agreement in category assignments using pairwise Cohen’s Kappa. Section˜D.1 includes further details on the classification process and on the agreement of GPT-4o with manual human annotations.
BLEU The BLEU surface similarity score (Papineni et al., 2002) between the generated question and the true question.
BertScore (Zhang et al., 2020) provides a cosine similarity score between the generated question and the true question based on contextual embeddings extracted from RoBERTa (Liu et al., 2019).
QA accuracy We introduce a downstream evaluation method with the following logic: the closer the generated question is to the true question, the easier it should be to select the correct answer for the true question when replaced with the generated question. We consider a question as a valid reconstruction of the true question if and only if, given this question, the text, and the four answers for the true question, the model selects the correct answer. We implement this evaluation using RoBERTa (Liu et al., 2019) fine-tuned for multiple choice QA on RACE (Lai et al., 2017), on the 86% of OneStopQA questions that RoBERTa answers correctly. We chose this model because OneStopQA is not in its training data, thus avoiding possible data contamination with newer models.
Baselines
To facilitate model performance interpretation, we introduce the following baseline questions:
Incorrect human questions The two additional human-composed questions for each text available in OneStop, one for a different critical span and one for the same critical span as the true question.
Arbitrary question from an LLM We generate an arbitrary question for the text with GPT-4o, using the prompt in Section˜D.2.
Question from a text-only question decoding model An additional question is generated by a Llama 3.1 model trained to decode the correct question without eye movement data.
6 Results
6.1 Question Selection
Table˜2 presents the aggregated test accuracy across the 10 data splits for the three-way ‘All’ selection, as well as a breakdown into questions with Different Spans, and questions with the Same Span evaluations. A division of the results into the New Text, New Participant, and New Text & Participant regimes is reported in Tables˜4, 5 and 6 in Appendix˜E. Validation set results are reported in Tables˜7, 8, 9 and 10 in Appendix˜E.
Table 2: Question selection test accuracy aggregated across 10 cross-validation splits, with 95% confidence intervals. A majority question choice is identical to chance because the data is balanced across questions. As the data is not i.i.d. (each participant reads multiple paragraphs, and each paragraph is read by multiple participants), we test for differences between models using a linear mixed effects model with random intercepts and slopes for participants and paragraphs: $is\_correct\sim model+(model\mid participant)+(model\mid paragraph)$. The accuracy of all the models is compared to chance performance, where significant gains over this baseline are marked with ‘*’ $p<0.05$, ‘**’ $p<0.01$, and ‘***’ $p<0.001$. The best performing model in each evaluation regime is marked in bold. Significant drops compared to the best model are marked with ‘+’ $p<0.05$, ‘++’ $p<0.01$ and ‘+++’ $p<0.001$.
Model Type Model All
Different Spans
Same Span
Discriminative Chance / Majority Baseline $33.0\pm 0.4_{\text{+++}}$ $55.3\pm 0.4_{\text{+++}}$ $49.9\pm 0.4_{\text{+++}}$
Question Similarity to RT-Weighted Passage $33.4\pm 0.3_{\text{+++}}$ $55.1\pm 3.7_{\text{+++}}$ $50.0\pm 0.4_{\text{+++}}$
RT Similarity to Question-Word Similarities $34.2\pm 0.4^{*}_{\text{+++}}$ $54.8\pm 1.4_{\text{+++}}$ $51.1\pm 0.5_{\text{+++}}$
Haller RNN (Haller et al., 2022) $41.8\pm 0.6^{***}_{\text{+++}}$ $65.6\pm 0.5^{***}_{\text{+++}}$ $52.1\pm 0.5^{**}_{\text{+++}}$
RoBERTEye-Fixations (Shubi et al., 2024) $\mathbf{49.3}\pm 0.3^{***}$ $\mathbf{70.9}\pm 0.5^{***}$ $\mathbf{57.3}\pm 0.5^{***}$
Generative DalEye-LLaVA $33.5\pm 0.3_{\text{+++}}$ $57.0\pm 0.7^{**}_{\text{+++}}$ $49.4\pm 0.4_{\text{+++}}$
DalEye-Llama $39.2\pm 0.9^{***}_{\text{+++}}$ $61.4\pm 0.9^{***}_{\text{+++}}$ $52.5\pm 0.4^{***}_{\text{+++}}$
Both Reading-Time Informed Embedding Similarity baseline models perform at chance level across all three evaluation regimes. A possible reason for this outcome lies in the inherent noise in the distribution of reading times of a single participant over a single paragraph, which deems heuristic methods that rely only on this distribution to be ineffective. Haller RNN and RoBERTEye-Fixations perform well above chance, and outperform the generative models on all the evaluations. In the generative models category, DalEye-Llama outperforms DalEye-LLaVA, where the latter is at chance level. Table˜11 in Section˜E.1.1 suggests that the alternative prompt and eye movements representations for DalEye-Llama weaken its performance. The advantage of the discriminative models over the generative models is not surprising, as differently from the discriminative models, the generative models are not explicitly trained on the question selection task.
RoBERTEye-Fixations achieves the highest accuracy across all three evaluation regimes. In Table˜4 of Appendix˜E, we observe that this performance advantage is consistent across the New Text, New Participant, and New Text & Participant evaluations. Thus, the model is able to generalize not only to new participants but importantly also to new texts. Notably, it is also the only model that is well above chance in the Same Span regime with $57.3\%$ accuracy. On the one hand, this confirms that distinguishing between questions that ask about the same portion of the paragraph is indeed much more challenging than distinguishing between questions over different parts of the paragraph. At the same time, the above change performance of RoBERTEye-Fixations on the Same Span evaluation speaks to the fine-grained nature of the information that can be extracted about the reader’s information seeking goals from their eye movements alone.
6.2 Question Reconstruction
Based on the question selection results in Table˜2, we focus on the DalEye-Llama model in the evaluation of question reconstruction quality. Figure˜4 presents the evaluations in the New Participant and the New Text regimes. Table˜12 in Appendix˜E presents these results in numerical form, and additionally includes the New Text & Participant regime, which yields very similar results to the New Text regime. Figure˜7 in appendix˜E presents an example of DalEye-Llama outputs for the reconstruction task. Table˜13 in Appendix˜E presents reconstruction results for the alternative task prompt for DalEye-Llama, which yields similar results, and the alternative eye movement representations, which yield inferior results to those presented below.
Figure 4: Question reconstruction evaluations of DalEye-Llama for (1) the identity of the generated question word, (2) UIUC semantic category of the question, (3) BLEU score, (4) BERTScore, (5) downstream QA accuracy based on the answer selection of a multiple-choice QA model. DalEye-Llama performance is benchmarked against two human-composed questions for a different and the same critical span, an arbitrary question generated with GPT-4o, and a question from a text-only Llama 3.1 model trained on the question decoding task using only the text. Presented are means with 95% confidence intervals (bootstrapped, $n=1000$).
Importantly, in the New Participant evaluation, DalEye-Llama generated questions outperform all the baseline questions on all five metrics. This suggests that DalEye-Llama generalizes relatively well to new participants when the text appeared in the training set. However, its performance drops both in absolute terms and relative to the baselines in the more challenging regime involving new texts. Notably, in this regime the incorrect human question - same critical span has the highest BLEU, BERTScore, and QA accuracy. This is likely the result of lexical overlap stemming from sharing the critical span with the true question. However, this span-sharing constraint may hinder this baseline in the Question word and Question category evaluations, as sharing a span tends to yield questions that differ in these regards. The human question over a different span, which tends to perform best in these evaluations, is not subject to this constraint.
7 Related Work
We build on two primary lines of work: (i) psycholinguistic investigation of goal driven reading; and (ii) predictive models of cognitive state from eye movements in reading.
7.1 Analyses of Goal Based Reading
While most prior work in the psychology of reading and psycholinguistics focuses on ordinary reading for general comprehension Rayner et al. (2012), goal or task based reading has been acknowledged as a key frontier in this area (Radach & Kennedy, 2004). Several studies found differences in eye movements between ordinary reading and tasks that define a reading manner: skimming, speed reading and proofreading (Just et al., 1982; Kaakinen & Hyönä, 2010; Schotter et al., 2014; Strukelj & Niehorster, 2018; Chen et al., 2023). Additional studies have found differences between ordinary reading and tasks of target word search (Rayner & Raney, 1996; Rayner & Fischer, 1996), word verification (Radach et al., 2008), and topic search (White et al., 2015). Prior work also analyzed eye movements during human linguistic tasks, such as named entity annotation (Tomanek et al., 2010; Tokunaga et al., 2017). Differences in reading patterns were also found when readers adopt different perspectives on a given text (Kaakinen et al., 2002; Kaakinen & Hyönä, 2008), or given different learning goals (Rothkopf & Billington, 1979). Overall, these works consistently found a systematic influence of the reading task on the resulting reading patterns.
Our work builds most directly on Kaakinen et al. (2015), Malmaud et al. (2020), Hahn & Keller (2023), and Shubi & Berzak (2023), who compared eye movements in ordinary reading and open-ended information seeking. These studies showed task conditioning effects on reading patterns, especially as a function of the task-relevance of the textual information. Similarly to these studies, we formulate the information seeking task as a reading comprehension question specific to the text. We further leverage the previously observed task conditioning effects for automatic discrimination between different tasks over the same text.
7.2 Decoding Cognitive State from Eye Movements
Multimodal modeling of eye movements and text has been a growing area of research in recent years (Reich et al., 2025), with applications ranging from improving NLP (e.g. Klerke et al., 2016; Mishra et al., 2016; Barrett & Hollenstein, 2020; Sood et al., 2020; Deng et al., 2024; López-Cardona et al., 2025) to prediction of eye movement trajectories (e.g. Hollenstein et al., 2021a; Bolliger et al., 2023; Deng et al., 2023). Within this broad line of research, several studies focused on cognitive state decoding: prediction of language-related aspects of the reader and their cognitive state from their eye movements while reading. These include studies on prediction of the reader’s linguistic background (Berzak et al., 2017; Reich et al., 2022; Skerath et al., 2023), language proficiency (Berzak et al., 2018), reading comprehension (Ahn et al., 2020; Reich et al., 2022; Shubi et al., 2024), and subjective readability (Reich et al., 2022).
Most notably, two studies addressed decoding of the reader’s goals. Hollenstein et al. (2021b) classified ordinary reading for comprehension versus annotation of semantic relations in single sentences using the ZuCo corpus (Hollenstein et al., 2020; 2023). Shubi et al. (2025) classified ordinary reading versus information seeking using the OneStop dataset. Importantly, in both studies, the decoding tasks are procedural, where the goal is to distinguish between a small set of pre-defined manners of reading that are not specific to a particular text. In contrast, our study addresses a related but conceptually and practically different task, which is not procedural, but rather semantic, decoding reading tasks specific to the text, within the information seeking regime. Instead of a small number of categories that can apply to any text, we have hundreds of text-specific tasks that can be arbitrary in nature. See further discussion on the procedural versus semantic tasks distinction in Shubi & Berzak (2023). Furthermore, these and other studies on cognitive state decoding typically use discriminative, encoder-based models that are limited to classification tasks. Here, we also integrate eye movements into generative, decoder-based language models, and investigate both textual- and embedding-based methods for conditioning text generation on eye movements in reading.
8 Summary and Discussion
We introduce a new challenge of both scientific and practical value: decoding of arbitrary information seeking goals over specific texts from eye movements in reading. We tackle this challenge in two formulations, goal selection and goal reconstruction, using a number of evaluation frameworks and modeling approaches. The best performing model on the selection task is able to predict the correct question with a considerable degree of success, even when the candidate reading goals pertain to the same information in the text. Our results further suggest that although the reconstruction task is extremely challenging, meaningful progress can also be made in this open-ended regime. Overall, we find that despite the inherent noise in eye movement data, effective modeling can extract highly valuable information regarding specific information seeking goals at the challenging granularity level of eye movements of a single reader over a single paragraph.
Automatic decoding of information seeking goals in real time can pave the way for new applications with positive societal value. In the future, media outlets and providers of municipal and governmental services could better understand the information needs of users who access their websites and render information in a manner that better meets these needs. Automatic identification of information seeking goals can also be used for real-time assistance to special populations, like elderly people, when accessing critical information on the web. Future e-learning systems could assess students’ progress on various types of information-seeking tasks and monitor their information-seeking skills over time. Finally, one can imagine interactive reading interfaces that include real-time text personalization, such as simplification and suggestion of additional relevant information, according to the reader’s information seeking goals. The present study is a first step in enabling such applications. Ample room remains for further improvement of prediction accuracy via modeling innovations, as well as the collection and analysis of data for additional variants of information seeking regimes, corpora, reader populations, and languages.
9 Ethical Considerations
The OneStop Eye Movements dataset used in this work was collected by Berzak et al. (2025). The study was conducted under an institutional IRB protocol, and all the participants provided written consent before participation. The data is anonymized. Analyses and predictive modeling of task-based reading are among the primary use cases for which the data was collected.
While the current work is a scientific proof of concept and is not aimed at a particular user-facing application, one has to consider potential use cases and risks for the presented task. In particular, it is important to note that given the current level of accuracy of goal decoding, educational and assistive technologies that rely on this task may be unreliable and introduce biases for various individuals and groups, such as L2 readers and groups with reading and cognitive impairments. Additional data collection and analyses are needed to assess such biases.
Prior work has demonstrated that eye movements can be used for user identification (e.g. Bednarik et al., 2005; Jäger et al., 2020). We do not perform user identification in this study. We further emphasize that future reading goal decoding technologies must be used in real-world applications only with explicit consent from potential users to have their eye movements collected and analyzed for this purpose.
10 Reproducibility Statement
We describe the model training and selection procedure, the evaluation protocol, the hyperparameter search space for each model, and the hardware and software specifications in Sections˜5 and C. The OneStop dataset (Berzak et al., 2025) used in the experiments and described in Section˜2 is publicly available at https://osf.io/2prdq/. The code for the paper, which implements all the models, experimental procedures, and analyses, is available at: https://anonymous.4open.science/r/open-question-prediction/.
Appendix
Appendix A Discriminative Models
The figures and descriptions below, depicting model architectures, show the processing of a single candidate question. As mentioned in Section˜4, each candidate question is processed independently.
A.1 Haller RNN
Given a text (a paragraph in our case) $T$ and a candidate question $q\in Q_{T}$, the model suggested by Haller et al. (2022) and adapted by us performs the following steps:
  1. 1.
    Word Embeddings: LLM-based contextualized word embeddings, for $T$ and $q$, resulting in three subsequences: (1) $Z_{[CLS]}$, (2) $Z_{q_{t_{1}}},\dots,Z_{q_{t_{l}}}]$, and (3) $Z_{T_{t_{1}}},\dots,Z_{T_{t_{x}}}]$.
  2. 2.
    Fixation Sequence: Paragraph-only token-level embeddings are pooled to word-level embeddings, and ordered by the fixation sequence. Then, each embedding in the sequence is concatenated with the respective fixation features. The fixation-level features are:
    • Horizontal position of the fixation
    • Total gaze duration (sum of all fixations on the word)
    • Duration of first fixation
    • Duration of outgoing saccade
    • Horizontal distance of outgoing saccade
    • Vertical distance of outgoing saccade
    • Total distance of outgoing saccade
    • Duration of incoming saccade
    • Horizontal distance of incoming saccade
    • Vertical distance of incoming saccade
  3. 3.
    Projection: $Z_{[CLS]}$ and $Z_{q_{t_{1}}},\dots,Z_{q_{t_{l}}}]$ are projected to a higher dimension to fit the dimension of the concatenated embeddings from the previous step.
  4. 4.
    RNN Input Sequence: The processed embeddings are then concatenated one after the other to create the RNN input sequence.
  5. 5.
    Classifier: Finally, a classifier layer scores each candidate question.
Figure˜5(a) depicts the architecture adapted to our task.
A.2 RoBERTEye-Fixations
Formally, given a paragraph $T$ and a candidate question $q\in Q_{T}$, the model constructs two parallel representations:
  1. 1.
    Textual Representation ($Z_{W}$): The input follows the format $Z_{W}=[\texttt{CLS};T;\texttt{SEP};q;\texttt{SEP}]$.
  2. 2.
    Fixation-Level Eye Movement Representation ($Z_{E_{T}}$): The eye movement sequence is defined as $Z_{E_{T}}=[Z_{E_{f_{1}}},...,Z_{E_{f_{m}}}]$. The fixation representation is computed as:
    $Z_{E_{f_{i}}}=\text{FC}(E_{f_{i}})+\text{Emb}_{\text{pos}}(w_{i})+\text{Emb}_{\text{eye}}$
    Here, $E_{f_{i}}$ captures fixation properties (e.g., duration, position) for fixation $f_{i}$ on word $w_{i}$. The fully connected layer FC projects this feature vector into the word embedding space, $\text{Emb}_{\text{pos}}(w_{i})$ is the positional embedding of $w_{i}$ used to map each fixation to the corresponding word, and $\text{Emb}_{\text{eye}}$ is a learnable embedding marking the presence of eye movement information. The fixation-level features are described in Table˜3.
  3. 3.
    Fusion and Prediction: The combined sequence $Z_{E_{T}};\texttt{SEP}_{E};Z_{W}]$ is processed by the transformer encoder, and the final CLS token representation is passed through two fully connected layers to predict the question: $\hat{q}=\operatorname{argmax}\;\text{MLP}(\texttt{CLS})$.
Figure˜5(b) depicts the adapted architecture.
Table 3: Eye movement and word property features used by RoBERTEye-Fixations and DalEye-Llama. See Berzak et al. (2025) for further details.
Feature Name Description
Word-Level Eye Movement Features
IA_DWELL_TIME The sum of the duration across all fixations that fell in the current interest area
IA_DWELL_TIME_% Percentage of trial time spent on the current interest area.
IA_FIXATION_% Percentage of all fixations in a trial falling in the current interest area.
IA_FIXATION_COUNT Total number of fixations falling in the interest area.
IA_REGRESSION_IN_COUNT Number of times interest area was entered from a higher IA_ID.
IA_REGRESSION_OUT_FULL_COUNT Number of times interest area was exited to a lower IA_ID.
IA_RUN_COUNT Number of times the Interest Area was entered and left (runs).
IA_FIRST_FIX_PROGRESSIVE Checks whether the first fixation in the interest area is a first-pass fixation.
IA_FIRST_FIXATION_DURATION Duration of the first fixation event that was within the current interest area
IA_FIRST_FIXATION_VISITED_IA_COUNT Number of distinct interest areas visited before the first fixation on the current area.
IA_FIRST_RUN_DWELL_TIME Dwell time of the first run.
IA_FIRST_RUN_FIXATION_COUNT Number of all fixations in a trial falling in the first run of the current interest area.
IA_SKIP IA_SKIP = 1 if the area received no fixation in first-pass reading.
IA_REGRESSION_PATH_DURATION Total fixation duration on the current interest area until moving to a higher IA_ID.
IA_REGRESSION_OUT_COUNT Exits from current IA to lower IDs before fixating a higher ID.
IA_SELECTIVE_REGRESSION_
PATH_DURATION
Duration of fixations and refixations on the current
interest area until eyes move to a higher ID.
IA_LAST_FIXATION_DURATION Duration of the last fixation event that was within the current interest area.
IA_LAST_RUN_DWELL_TIME Dwell time of the last run.
IA_LAST_RUN_FIXATION_COUNT Number of fixations during the last run (sequence) within the current interest area.
IA_TOP Y coordinate of the top of the interest area.
IA_LEFT X coordinate of the left-most part of the interest area.
IA_FIRST_FIX_PROGRESSIVE Whether the first fixation in the interest area is progressive (left-to-right, first-pass).
normalized_Word_ID Position in the paragraph of the word interest area, normalized from zero to one.
PARAGRAPH_RT Reading time of the entire paragraph.
total_skip Binary indicator whether the word was fixated on.
Fixation-level Eye Movement Features
CURRENT_FIX_INDEX The position of the current fixation in the trial.
CURRENT_FIX_DURATION Duration of the current fixation.
CURRENT_FIX_PUPIL Average pupil size during the current fixation.
CURRENT_FIX_X X coordinate of the current fixation.
CURRENT_FIX_Y Y coordinate of the current fixation.
NEXT_FIX_ANGLE, PREVIOUS_FIX_ANGLE Angle between the horizontal and the line from the current to the next/previous fixation.
NEXT/PREVIOUS_FIX_DISTANCE Distance between current fixation and next/previous fixation.
NEXT_SAC_AMPLITUDE Amplitude of the following saccade in degrees of visual angle.
NEXT_SAC_ANGLE Angle between the horizontal plane and the direction of the next saccade.
NEXT_SAC_AVG_VELOCITY Average velocity of the next saccade.
NEXT_SAC_DURATION Duration of the next saccade in milliseconds.
NEXT_SAC_PEAK_VELOCITY Peak values of gaze velocity (in visual degrees per second) of the next saccade.
NEXT_FIX_INTEREST_AREA_INDEX Index of the interest area where the next fixation lands
Word Properties
gpt2_surprisal Difficulty predicted by GPT-2 model
wordfreq_frequency Frequency of word in language usage
word_length Number of characters in the word
start_of_line Indicates word is at start of line
end_of_line Indicates word is at end of line
is_content_word Is the word a content word (noun, verb, etc.)
left_dependents_count Count of dependents left of word
right_dependents_count Count of dependents right of word
distance_to_head Syntactic distance to head word
(a) Haller RNN
(b) RoBERTEye-Fixations
Figure 5: Visualization of the Haller RNN and RoBERTEye-Fixations architectures. LMLM stands for a language model, RNNRNN for a recurrent neural network. $FF_{f_{i}}$ stands for the fixation features and $w_{f_{i}}$ for the word corresponding to the $i$-th fixation respectively.
Appendix B Generative Models
B.1 DalEye-LLaVA
DalEye-LLaVA extends the LLaVA-OneVision architecture by replacing the vision encoder with a fixation encoder tailored to eye movement features. The model input is structured as an instruction-style prompt comprising three components: (1) a task description, (2) the paragraph text, and (3) the participant’s eye movement features, positionally aligned with the paragraph words. The prompt concludes with the ground-truth question followed by an <eos> token. The entire sequence is tokenized, and training uses teacher forcing, with the cross-entropy loss applied only to the question tokens.
The forward architecture mirrors LLaVA’s design, with three main stages, as depicted in Figure˜6:
  1. 1.
    Text encoding: The task prompt and paragraph tokens are embedded by the backbone language model.
  2. 2.
    Fixation encoding: Eye movement features (see Table˜3) are processed by a dedicated fixation encoder, which applies MLP layers across the feature dimension and 1D convolutions across the temporal dimension, producing a sequence of fixation embeddings.
  3. 3.
    Fusion: Fixation embeddings replace the <image> placeholder token in the input sequence. Word tokens are assigned positional IDs according to their order in the paragraph, while each fixation embedding inherits the positional ID of the corresponding fixated word.
Figure 6: DalEye-LLaVA architecture. Text tokens and eye movement features are fused via a fixation encoder and rotary positional embeddings before being passed to the LLM.
B.2 DalEye-Llama
In Section˜B.2.1 we detail the task prompt and eye-movement input representations used for question generation with DalEye-Llama. Variants of these prompts and inputs are described in Section˜B.2.2.
B.2.1 Input format
You will be given data from an eye-tracking for reading experiment in which participants first read a question about a paragraph, then read the paragraph, and finally answered the question.
Input: You will be provided with a paragraph and eye movements of a single participant over that paragraph.
Output: Your task is to generate the question that was presented to the participant prior to reading the paragraph.
Eye Movements Representation:
You will receive the eye movements data for the paragraph formatted as <FORMAT>
Instructions:
Output only the original question presented to the reader, matching it as best as you can. DO NOT include any additional commentary or explanation.
<PARAGRAPH>
<SCANPATH>
<PARAGRAPH> is the textual paragraph, for example: ‘‘The quick brown fox jumps...’’.
<SCANPATH> is the eye movements data which depends on the <FORMAT>.
B.2.2 Input and Task Prompt Variations
We experiment with three input representations that capture different granularities of eye movement information: (i) Fixation—level, (ii) Word-level, and (iii) a combination of both fixation- and word-level information, as detailed below:
  1. 1.
    Fixations: Fixation Duration + Next saccade direction
    • Format: a list of fixation-level features: [fixated word index, fixated word, fixation duration in ms, direction of next saccade (backward to earlier word / within word / forward to later word), words between current and next fixation]
    • Example: [[4, ‘‘fox’’, 220, backward], ...]
  2. 2.
    Words: Total Fixation Duration + Incoming/Outgoing Backward/Forward Saccades
    • Format: a list of word-level features: [word index, word, total fixation duration in ms, incoming forward saccades (from earlier word), incoming backward saccades (from later word), outgoing forward saccades (to later word), outgoing backward saccades (to earlier word)]
    • Example: [[4, ‘‘fox’’, 320, 2, 1, 3, 0], ...]
  3. 3.
    Words+Fixations
    • Format: two lists of word-level and fixation-level features: [fixated word index, fixated word, fixation duration in ms, direction of next saccade (backward / within / forward)] [word index, word, total fixation duration in ms, incoming forward saccades, incoming backward saccades, outgoing forward saccades, outgoing backward saccades]
    • Example: [[4, ‘‘fox’’, 220, backward], ...] [[4, ‘‘fox’’, 320, 2, 1, 3, 0], ...]
We further experiment with an alternative prompt that de-emphasizes the experimental setup and places more focus on the model’s generation task:
Task Description:
Your task is to generate the exact original question a reader had in mind before reading a given paragraph. The input data is (a/two) time series composed of (word-level/fixation-level/word-level and fixation-level) features.
Input Format:
You will receive the paragraph and eye movement data formatted as <FORMAT>
Expected Output:
Generate the exact original question provided to the reader, accurately inferred from the fixation patterns.
Instructions:
Your output must precisely match the original question presented to the reader.
Produce only the exact original question as your output.
<PARAGRAPH>
<SCANPATH>
Appendix C Model Training and Hyperparameters
We use a stringent evaluation protocol, in which paragraphs are assigned to training, validation, and test sets at the article level, such that all the paragraphs from the same article appear in the same portion of each data split.
C.1 Discriminative Models
Since the models we use were developed for different tasks and datasets, we conducted a hyperparameter search for each model. The search space for each model is described below. In all cases, it includes the optimal parameters reported in the work that introduced the model, extended to provide a fair comparison between models.
For all neural models, we train with learning rates of $\{0.00001,0.00003,0.0001,0.0002\}$, following Shubi et al. (2024). Additionally, for all models that make use of word embeddings, we include both frozen and unfrozen language model variants in the search space.
  • For Haller RNN, we search over LSTM hidden sizes of $\{10,40,70,140\}$, using one layer and a dropout rate of $0.1$. The model is trained with and without freezing language model parameters.
  • For RoBERTa-F, we search over dropout rates of $\{0.1,0.3,0.5\}$ for the eye movement projection layer. The model is trained with and without freezing language model parameters.
We train the deep-learning-based models for a maximum of 40 epochs, with early stopping after 8 epochs if no improvement in the validation error is observed. A single training epoch took roughly 5 minutes for RoBERTa-F, 10 minutes for Haller RNN, and 30 minutes for Llava and Llama. Each individual run was capped at 24 hours. Following Liu et al. (2019); Mosbach et al. (2021); Shubi et al. (2024), we use the AdamW optimizer (Loshchilov & Hutter, 2018) with a batch size of $16$. RoBERTa-F uses a linear warm-up ratio of $0.06$ and a weight decay of $0.1$. We standardize each eye movement feature using statistics computed on the training set, to zero mean unit variance.
Both reading-time informed embedding similarity baseline models use word embeddings from the RoBERTa-Large (Liu et al., 2019) language model.
C.2 Generative Models
We present two generative models, DalEye-Llama and DalEye-LLaVA, each fine-tuned with distinct methods and hyperparameters optimized for their specific architectures and objectives.
DalEye-Llama
This model is fine-tuned using Unsloth with the Meta-Llama-3.1-8B backbone loaded in 4-bit precision. Training employs Low-Rank Adaptation (LoRA) with rank $r=16$, scaling factor $\alpha=16$, and applies RS-LoRA regularization. LoRA targets transformer modules: q_proj, k_proj, v_proj, up_proj, down_proj, o_proj, and gate_proj. The model is trained for 2 epochs with a linear scheduler and warm-up of 10 steps, batch size of 1, gradient accumulation over 2 steps, AdamW-8bit optimizer with learning rate $1\times 10^{-4}$, and weight decay of 0.01.
DalEye-LLaVA
This model employs a teacher-forcing strategy. Specifically, loss computation is restricted exclusively to tokens corresponding to possible questions, contrasting the likelihood of generating the correct question against two distractors. The computed losses for each question candidate are inverted, serving as logits for softmax normalization and subsequently optimized via cross-entropy against the true label. The LLaVA backbone remains frozen, and training employs LoRA with dropout rate of 0.1, RS-LoRA, and rank $r=8$. The hyperparameter search includes learning rates of $1\times 10^{-5},3\times 10^{-5},1\times 10^{-4}$}, fixation embedding hidden sizes of 512,1024{512,1024}, and early stopping after 8 epochs without validation improvement.
All neural network-based models were trained using the PyTorch Lightning (Falcon & team, 2024) library on NVIDIA A100-40GB and L40S-48GB GPUs.
We utilize the Hugging Face implementation of LLaVA-OneVision, specifically LLaVA-hf/LLaVA-onevision-qwen2-0.5b-si-hf. Additionally, we employ the gpt-4o-2024-08-06 version of GPT-4o. We use Unsloth (version 2025.9.7) (Daniel Han & team, 2023) to fine-tune the Llama model. For the QA model we use LIAMF-USP/roberta-large-finetuned-race.
There are roughly 355M trainable parameters for RoBERTa-F, reduced to 3M when the RoBERTa backbone is frozen. Haller RNN consists of 357M trainable parameters, or 2.4M when the backbone is frozen. DalEye-LLaVA has 505M parameters, out of which roughly 12M were unfrozen. DaLEye-Llama has 8.1B parameters, out of which roughly 42M were unfrozen.
Appendix D Question Reconstruction
D.1 Question Annotation into UIUC Categories
This section provides additional information regarding the question annotation into UIUC categories. Specifically, in Section˜D.1.1 we include the full prompt given to the model, in Section˜D.1.2 examples of questions and their corresponding annotations, and in Section˜D.1.3 agreement with human annotators.
D.1.1 Full Prompt
You are an expert at classifying questions into the standard UIUC
**main question categories**.
The main categories include:
- ABBR: Abbreviation
For example:
  - What does the abbreviation AIDS stand for?
  - What is the abbreviation for micro?
  - What is the abbreviation of the company name 'General Motors'?
  - What does G.M.T. stand for?
- ENTITY: Entity
For example:
  - What kind of animal is Babar?
  - What killed Bob Marley?
  - What is a fear of weakness?
  - Where does your hair grow the fastest?
- DESCRIPTION: Description
For example:
  - What do Mormons believe?
  - What is the history of skateboarding?
  - What is the difference between a generator and an alternator?
  - Where do rocks come from?
- MANNER: Manner
For example:
  - How do I get another city's newspaper?
  - How do you solve "Rubik's Cube"?
  - How do you look up criminal records on the Internet?
  - How do you find oxidation numbers?
- REASON: Reason
For example:
  - What is the purpose of a car bra?
  - What makes a tornado turn?
  - What causes the redness in your cheeks when you blush?
  - Why do horseshoes bring luck?
- DEFINITION: Definition
For example:
  - What is a dental root canal?
  - What is the contents of proposition 98?
  - Hazmat stands for what?
  - What does the name Billie mean?
- HUMAN: People or groups
For example:
  - Who invented baseball?
  - What stereo manufacturer is 'Slightly ahead of its time'?
  - Who played the original Charlie's Angels?
  - What company's logo is a 'W' in a circle?
- LOCATION: Geographic locations
For example:
  - Where is the highest point in Japan?
  - What European city do Nicois live in?
  - What is Answers.com 's address?
  - What U.S. state borders Illinois to the north?
- NUMERIC: Numbers, quantities, and dates
For example:
  - What is the temperature for baking Peachy Oat Muffins?
  - How many colleges are in Wyoming?
  - What is the average temperature in the Arctic?
  - What is the speed of light?
Return an array of tuples in the format:
[ (0, "HUMAN"), (1, "DESCRIPTION"), (2, "NUMERIC"), (3, "DESCRIPTION"), ]
If unsure, choose the closest matching category.
D.1.2 Categorization Examples
For each UIUC question category we present an example question sourced from OneStopQA Berzak et al. (2020):
  • REASON - Why does Myslajek mention Russia, Lithuania and Belarus?
  • NUMERIC - Approximately how many taxi drivers are there in the UK?
  • MANNER - How does Myslajek react to what he sees between the two paw prints?
  • LOCATION - Where was wolf-hunting banned in 1995?
  • ENTITY - Which of the following will be featured at Pestival 2013?
  • HUMAN - Who threw the bottle into the Baltic Sea?
  • DESCRIPTION - What does Angella think of the state of the sea today?
D.1.3 Agreement with Human Annotators
To evaluate the quality of the GPT-4o question category classifications, we randomly sampled 100 questions. A human annotator (one of the paper’s authors) manually labeled them with a UIUC category. We find an 86% agreement (0.794 Cohen’s kappa) between the human annotations and the model classifications.
D.2 Arbitrary Question Generation
Below is the prompt used for generating questions using Llama and GPT-4o.
Task Description:
Your task is to generate a question a reader had in mind before reading a given paragraph. The input data is the paragraph itself in a standard textual format.
Input Format:
You will receive a paragraph in plain text format:
[Paragraph text]
Expected Output:
Generate the exact original question provided to the reader, accurately inferred from the paragraph content. Identify key themes, concepts, or statements in the paragraph that strongly indicate the question that initially motivated the reading.
Instructions:
Your output must precisely match the original question presented to the reader.
Focus specifically on central concepts, themes, or statements within the paragraph.
Produce only the exact original question as your output.
Appendix E Additional Results
E.1 Question Selection
The tables below present a breakdown of the test and validation results by the All Spans, Different Spans and Same Spans tasks, and by New Item, New Participant and New Item & Participant evaluation regimes aggregated across 10 cross-validation splits.
Model performance is compared to the Majority class baseline using a linear mixed effects model. In R notation: $is\_correct\sim model+(model\mid participant)+(model\mid paragraph)$. Significant gains over this baseline are marked with ’*’ $p<0.05$, ’**’ $p<0.01$ and ’***’ $p<0.001$ in superscript. The best performing model in each evaluation regime is marked in bold. Significant drops compared to the best model are marked in subscript with ’+’.
Table 4: Test accuracy results for the All Spans task by evaluation regime.
Model New Item New Participant New Item & Participant
Chance / Majority Baseline $2.7\pm 0.6_{\text{+++}} 33$$2\pm 0.5_{\text{+++}} 33.$$3.3\pm 0.9_{\text{+++}}$
Question Similarity to RT-Weighted Passage $3.4\pm 0.4_{\text{+++}} 33$$3.3\pm 0.4_{\text{+++}}$ $34.9\pm 1.4_{\text{+++}}$
Reading Times Similarity to Question Paragraph Similarities $34.1\pm 0.6^{*}_{\text{+++}}$ $34.3\pm 0.6_{\text{+++}}$ $34.2\pm 1.2_{\text{+++}}$
Haller RNN (Haller et al., 2022) $40.4\pm 0.5^{***}_{\text{+++}}$ $43.3\pm 1.1^{***}_{\text{+++}}$ $40.8\pm 1.4^{***}_{\text{++}}$
RoBERTEye-Fixations (Shubi et al., 2024) $\mathbf{49.0}\pm 0.8^{***}$ $\mathbf{49.9}\pm 0.7^{***}$ $\mathbf{47.8}\pm 1.5^{***}$
DalEye-LLaVA $3.6\pm 0.2_{\text{+++}} 33$$3.2\pm 0.3_{\text{+++}}$ $34.6\pm 1.1_{\text{+++}}$
DalEye-Llama $35.4\pm 0.4^{***}_{\text{+++}}$ $43.1\pm 1.6^{***}_{\text{+++}}$ $39.3\pm 1.0^{**}_{\text{+++}}$
Table 5: Test accuracy results for the Different Spans task variation by evaluation regime.
Model
New Item
Different Spans
New Participant
Different Spans
New Item & Participant
Different Spans
Chance / Majority Baseline $5.2\pm 0.7_{\text{+++}} 55$$5.7\pm 0.5_{\text{+++}}$ $53.5\pm 1.5_{\text{+++}}$
Question Similarity to RT-Weighted Passage $55.1\pm 3.7_{\text{+++}}$ $54.9\pm 3.7_{\text{+++}}$ $57.8\pm 3.6_{\text{+++}}$
Reading Times Similarity to Question Paragraph Similarities $54.8\pm 1.4_{\text{+++}}$ $4.8\pm 1.4_{\text{+++}} 55$$5.1\pm 1.7_{\text{+++}}$
Haller RNN (Haller et al., 2022) $4.9\pm 0.6^{***}_{\text{+++}} 66$$2\pm 0.9^{***}_{\text{+++}} 66.$$6.4\pm 1.2^{***}$
RoBERTEye-Fixations (Shubi et al., 2024) $\mathbf{71.2}\pm 0.9^{***}$ $\mathbf{70.7}\pm 0.7^{***}$ $\mathbf{69.1}\pm 1.1^{***}$
DalEye-LLaVA $8.1\pm 1.2^{***}_{\text{+++}} 55$$5.7\pm 0.5_{\text{+++}}$ $58.7\pm 1.8^{*}_{\text{+++}}$
DalEye-Llama $57.4\pm 1.0^{*}_{\text{+++}}$ $65.7\pm 1.4^{***}_{\text{+++}}$ $59.9\pm 0.9^{**}_{\text{+++}}$
Table 6: Test accuracy results for the Same Spans task variation by evaluation regime.
Model
New Item
Same Span
New Participant
Same Span
New Item & Participant
Same Span
Chance / Majority Baseline $49.4\pm 0.5_{\text{{+++}}}$ $50.2\pm 0.6_{\text{{+++}}}$ $51.4\pm 1.8$
Question Similarity to RT-Weighted Passage $50.3\pm 0.5_{\text{{+++}}}$ $49.8\pm 0.8_{\text{{+++}}}$ $49.7\pm 2.3_{\text{{++}}}$
RT Similarity to Question-Word Similarities $51.0\pm 0.7_{\text{{+++}}}$ $51.0\pm 0.5_{\text{{+++}}}$ $53.1\pm 1.4$
Haller RNN (Haller et al., 2022) $51.7\pm 0.5^{*}_{\text{{+++}}}$ $52.6\pm 1.0^{*}_{\text{{+++}}}$ $50.3\pm 2.0_{\text{{+}}}$
RoBERTEye-Fixations (Shubi et al., 2024) $\mathbf{56.5}\pm 0.8^{***}$ $\mathbf{58.2}\pm 0.7^{***}$ $\mathbf{56.7}\pm 1.6$
DalEye-LLaVA $49.6\pm 0.4_{\text{{+++}}}$ $49.0\pm 0.4_{\text{{+++}}}$ $51.5\pm 1.6_{\text{{+}}}$
DalEye-Llama $50.4\pm 0.3_{\text{{+++}}}$ $4.2\pm 0.7^{***}_{\text{{+++}}} 55$$5.0\pm 1.5$
Table 7: Validation accuracy results for each of the tasks.
Model All Spans
Diff Span
Same Span
Chance / Majority Baseline $33.1\pm 0.3_{\text{+++}}$ $55.5\pm 0.4_{\text{+++}}$ $50.1\pm 0.3_{\text{+++}}$
Question Similarity to RT-Weighted Passage $33.0\pm 0.4_{\text{+++}}$ $54.6\pm 3.8_{\text{+++}}$ $50.0\pm 0.5_{\text{+++}}$
Reading times similarity to question-word similarities $34.1\pm 0.3_{\text{+++}}$ $54.3\pm 1.3_{\text{+++}}$ $51.1\pm 0.5_{\text{+++}}$
Haller RNN (Haller et al., 2022) $44.2\pm 0.3^{***}_{\text{+++}}$ $66.4\pm 0.4^{***}_{\text{+++}}$ $53.3\pm 0.4^{***}_{\text{+++}}$
RoBERTEye-Fixations (Shubi et al., 2024) $\textbf{50.9}\pm 0\textbf{.3}^{***}$ $\textbf{71.5}\pm\textbf{0.4}^{***}$ $\textbf{58.5}\pm\textbf{0.5}^{***}$
DalEye-LLaVA $34.8\pm 0.3^{**}_{\text{+++}}$ $56.8\pm 0.7^{*}_{\text{+++}}$ $50.8\pm 0.4_{\text{+++}}$
DalEye-Llama $38.4\pm 0.4^{***}_{\text{+++}}$ $60.5\pm 0.6^{***}_{\text{+++}}$ $51.7\pm 0.4^{*}_{\text{+++}}$
Table 8: Validation accuracy results for the All Spans task by evaluation regime.
Model New Item New Participant New Item & Participant
Chance / Majority Baseline $3.3\pm 0.5_{\text{+++}} 33$$3.0\pm 0.6_{\text{+++}}$ $31.7\pm 1.2_{\text{+++}}$
Question Similarity to RT-Weighted Passage $2.6\pm 0.3_{\text{+++}} 33$$3.2\pm 0.7_{\text{+++}}$ $34.7\pm 1.6_{\text{+++}}$
RT Similarity to Question-Word Similarities $34.1\pm 0.6_{\text{+++}}$ $34.2\pm 0.6_{\text{+++}}$ $34.3\pm 1.3_{\text{+++}}$
Haller RNN (Haller et al., 2022) $42.9\pm 0.3^{***}_{\text{+++}}$ $45.7\pm 0.8^{***}_{\text{+++}}$ $42.4\pm 1.1^{***}_{\text{+++}}$
RoBERTEye-Fixations (Shubi et al., 2024) $\textbf{50.6}\pm\textbf{0.8}^{***}$ $\textbf{51.3}\pm\textbf{0.7}^{***}$ $\textbf{50.7}\pm\textbf{1.5}^{***}$
DalEye-LLaVA $34.3\pm 0.3_{\text{+++}}$ $34.8\pm 0.6^{*}_{\text{+++}}$ $38.5\pm 1.5^{**}_{\text{+++}}$
DalEye-Llama $35.8\pm 0.4^{**}_{\text{+++}}$ $40.8\pm 0.7^{***}_{\text{+++}}$ $39.1\pm 1.1^{***}_{\text{+++}}$
Table 9: Validation accuracy results for the Different Spans task variation by evaluation regime.
Model
New Item
Different Spans
New Participant
Different Spans
New Item & Participant
Different Spans
Chance / Majority Baseline $5.8\pm 0.6_{\text{+++}} 55$$5.4\pm 0.6_{\text{+++}}$ $53.2\pm 0.9_{\text{+++}}$
Question Similarity to RT-Weighted Passage $54.6\pm 3.8_{\text{+++}}$ $54.4\pm 3.9_{\text{+++}}$ $57.3\pm 3.8_{\text{+++}}$
RT Similarity to Question-Word Similarities $54.3\pm 1.4_{\text{+++}}$ $54.4\pm 1.2_{\text{+++}}$ $53.8\pm 2.3_{\text{+++}}$
Haller RNN (Haller et al., 2022) $65.1\pm 0.8^{***}_{\text{+++}}$ $67.8\pm 1.0^{***}_{\text{+++}}$ $65.3\pm 0.7^{***}_{\text{+}}$
RoBERTEye-Fixations (Shubi et al., 2024) $\textbf{71.6}\pm\textbf{0.6}^{***}$ $\textbf{71.6}\pm\textbf{0.7}^{***}$ $\textbf{70.4}\pm\textbf{1.9}^{***}$
DalEye-LLaVA $56.6\pm 1.1_{\text{+++}}$ $56.4\pm 0.8_{\text{+++}}$ $60.8\pm 1.9^{***}_{\text{+++}}$
DalEye-Llama $58.5\pm 1.1^{*}_{\text{+++}}$ $62.6\pm 0.9^{***}_{\text{+++}}$ $61.2\pm 0.9^{***}_{\text{+++}}$
Table 10: Validation accuracy results for the Same Spans task variation by evaluation regime.
Model
New Item
Same Span
New Participant
Same Span
New Item & Participant
Same Span
Chance / Majority Baseline $50.1\pm 0.4_{\text{+++}}$ $50.0\pm 0.6_{\text{+++}}$ $50.7\pm 2.0_{\text{++}}$
Question Similarity to RT-Weighted Passage $49.3\pm 0.6_{\text{+++}}$ $50.6\pm 0.8_{\text{+++}}$ $49.9\pm 2.2_{\text{++}}$
RT Similarity to Question-Word Similarities $50.9\pm 0.6_{\text{+++}}$ $51.0\pm 0.6_{\text{+++}}$ $53.0\pm 1.5_{\text{+}}$
Haller RNN (Haller et al., 2022) $53.2\pm 0.4^{**}_{\text{+++}}$ $53.4\pm 0.7^{**}_{\text{+++}}$ $52.2\pm 1.7_{\text{+}}$
RoBERTEye-Fixations (Shubi et al., 2024) $\textbf{57.3}\pm\textbf{0.9}^{***}$ $\textbf{59.7}\pm\textbf{0.5}^{***}$ $\textbf{58.6}\pm\textbf{1.9}^{**}$
DalEye-LLaVA $51.0\pm 0.3_{\text{+++}}$ $50.7\pm 0.9_{\text{+++}}$ $49.8\pm 1.5_{\text{+++}}$
DalEye-Llama $50.9\pm 0.4_{\text{+++}}$ $52.5\pm 0.6^{*}_{\text{+++}}$ $53.2\pm 1.2_{\text{+}}$
E.1.1 DalEye-Llama Input Variations
In Table˜11 we present the results of using either one of the two proposed task prompts and three input representations - fixation-level, word-level, and a combined fixation- and word-level representation (see Section˜B.2.1 for the full inputs). Across all three task variations, Task prompt #1 with fixation-level input yields the best accuracy. Both the alternative task prompt and the word-level/combined inputs underperform this setting, where the choice of representation matters more than the task prompt.
Table 11: Test accuracy results for each of the tasks and for each of the input variations.
Task Prompt Representation All Spans
Diff Span
Same Span
Chance / Majority Baseline None $33.0\pm 0.4_{\text{+++}}$ $55.3\pm 0.4_{\text{+++}}$ $49.9\pm 0.4_{\text{+++}}$
Task prompt #1 Fixation-level $\textbf{39.2}\pm\textbf{0.9}^{***}$ $\textbf{61.4}\pm\textbf{0.9}^{***}$ $\textbf{52.5}\pm\textbf{0.4}^{***}$
Word-level $33.9\pm 0.5_{\text{+++}}$ $55.5\pm 0.6_{\text{+++}}$ $50.4\pm 0.3_{\text{++}}$
Both fixation- and word-level $34.7\pm 0.6^{***}_{\text{+++}}$ $56.2\pm 0.8_{\text{+++}}$ $50.5\pm 0.4_{\text{++}}$
Task prompt #2 Fixation-level $37.2\pm 0.6^{***}_{\text{+++}}$ $60.3\pm 0.6^{***}$ $51.0\pm 0.3_{\text{+}}$
Word-level $33.4\pm 0.2_{\text{+++}}$ $55.0\pm 0.6_{\text{+++}}$ $50.2\pm 0.2_{\text{+++}}$
Both fixation- and word-level $34.4\pm 0.7^{**}_{\text{+++}}$ $55.5\pm 1.0_{\text{+++}}$ $50.5\pm 0.5_{\text{++}}$
E.2 Question Reconstruction
Example generations for DalEye-Llama are presented in Figure˜7. Question reconstruction evaluations of the DalEye-Llama model for (1) the identity of the generated question word, (2) the semantic category of the question (Li & Roth, 2002), (3) similarity using BERTScore, and (4) downstream QA accuracy based on the answer selection of a multiple-choice question answering model. DalEye-Llama performance is benchmarked against two types of human composed questions (for a different critical span and for the same span) and against arbitrary questions generated with GPT-4o and Llama 3.1 in Table˜12 (same as Figure˜4, in numerical format). Presented are means with 95% confidence intervals (bootstrapped, $n=1000$) for three generalization levels to new participants and new readers. The highest score in each combination of evaluation and generalization level is marked in bold. In Table˜13 we present the results of DalEye-Llama compared to the word-level and combined word- and fixation-level input representations, and for the task prompt variation.
Paragraph: Lemnos has wild beaches, where you can swim and sunbathe almost alone, a small nightlife scene and many cultural sites. Lemnos is the eighth largest island in Greece so it will have to pay the first round of tax increases in autumn 2015. But Lemnos is far less wealthy than many smaller islands. It has just over 3,000 beds for visitors – Rhodes, for example, has tens of thousands of beds. "We have been suffering economically in recent years and now we will suffer more," said Lemnos Mayor, Dimitris Marinakis.
True Question New Item Generated Questions New Participant Generated Questions What is among the primary attractions of Lemnos? (A) What is one thing to note about the eponymous village? (A) What is among the primary attractions of Lemnos? (B) Who is quoted as saying that hotels benefit from the increase in the number of tourists in the area? (B) What will happen in spring 2013? (C) What does the quote convey about the current economic situation in Greece? (C) Why does Lemnos have to pay the first round of tax increases? (D) What is true of the other Greek islands? (D) What is one reason why the wealthy tend to spend more time on the islands?
Figure 7: An example showing a paragraph, the corresponding ground truth question, and four DalEye-Llama generated questions from the New Item Regime and four from the New Participant Regime. Each generated question reflects a different outcome from the QA model, as indicated by the answer choice (A–D) selected by the model. The possible answers, structured according to the STARC annotation guidelines (Berzak et al., 2020), to the ground truth question were: A) (correct answer) A large number of cultural destinations. B) (miscomprehension of the critical span) Beaches constantly full of locals and tourists. C) (incorrect, related to a different span) 3,000 luxury hotels. D) (no support in the passage) Highly-regarded restaurants.
Table 12: Test evaluations of the DalEye-Llama model benchmarked against two types of human composed questions (for a different critical span and for the same span) and against arbitrary questions generated with GPT-4o and Llama 3.1.
New Participant New Item
New Participant & Item
Question word (Kappa)
[…]
Incorrect human question - same critical span $087\pm 0.014 0.0$$068\pm 0.013$ $\mathbf{0.107\pm 0.040}$
GPT-4o arbitrary question $020\pm 0.008 0.0$$25\pm 0.007 0.0$$016\pm 0.019$
Llama 3.1 arbitrary question $346\pm 0.011 0.0$$12\pm 0.005 0.0$$003\pm 0.013$
DalEye-Llama question $\mathbf{0.478\pm 0.011}$ $069\pm 0.008 0.0$$052\pm 0.022$
Question category (Kappa)
Incorrect human question - different critical span $0.097\pm 0.009$ $\mathbf{0.107\pm 0.009}$ $\mathbf{0.133\pm 0.028}$
Incorrect human question - same critical span $063\pm 0.014 0.0$$046\pm 0.013$ $0.104\pm 0.040$
GPT-4o arbitrary question $005\pm 0.006 0.0$$18\pm 0.006 0.0$$022\pm 0.019$
Llama 3.1 arbitrary question $250\pm 0.011 0.0$$28\pm 0.009 0.0$$024\pm 0.026$
DalEye-Llama question $\mathbf{0.299\pm 0.011}$ $074\pm 0.010 0.0$$045\pm 0.028$
BLEU
Incorrect human question - different critical span $0.291\pm 0.002$ $0.291\pm 0.002$ $0.286\pm 0.007$
Incorrect human question - same critical span $0.384\pm 0.004$ $\mathbf{0.380\pm 0.004}$ $\mathbf{0.379\pm 0.011}$
GPT-4o arbitrary question $0.212\pm 0.002$ $0.217\pm 0.002$ $0.210\pm 0.006$
Llama 3.1 arbitrary question $0.523\pm 0.005$ $0.299\pm 0.003$ $0.309\pm 0.007$
DalEye-Llama question $\mathbf{0.602\pm 0.006}$ $0.315\pm 0.003$ $0.314\pm 0.007$
BERTScore
Incorrect human question - different critical span $0.604\pm 0.001$ $0.603\pm 0.001$ $0.601\pm 0.003$
Incorrect human question - same critical span $0.651\pm 0.002$ $\mathbf{0.650\pm 0.002}$ $\mathbf{0.654\pm 0.006}$
GPT-4o arbitrary question $0.413\pm 0.005$ $0.398\pm 0.005$ $0.313\pm 0.013$
Llama 3.1 arbitrary question $0.723\pm 0.003$ $0.617\pm 0.001$ $0.617\pm 0.003$
DalEye-Llama question $\mathbf{0.774\pm 0.003}$ $0.631\pm 0.002$ $0.628\pm 0.004$
QA accuracy
Incorrect human question - different critical span $49.7\pm 0.9$ $49.2\pm 0.8$ $48.7\pm 2.3$
Incorrect human question - same critical span $68.1\pm 1.1$ $\mathbf{67.7\pm 1.1}$ $66.3\pm 3.0$
[…]
Llama 3.1 arbitrary question $68.6\pm 0.8$ $60.6\pm 0.8$ $62.0\pm 2.3$
DalEye-Llama question $\mathbf{76.3\pm 0.8}$ $65.1\pm 0.8$ $65.2\pm 2.2$
Table 13: Test evaluations of the DalEye-Llama model benchmarked against the word-level and combined word- and fixation-level input representations, and for the task prompt variation.
New Participant New Item
New Participant & Item
Question word (Kappa)
Task prompt #1 Fixation-level $\mathbf{0.478\pm 0.011}$ $\mathbf{0.069\pm 0.008}$ $\mathbf{0.052\pm 0.022}$
Word-level $047\pm 0.004 0.0$$40\pm 0.005 0.0$$029\pm 0.012$
Both fixation- and word-level $054\pm 0.005 0.0$$40\pm 0.005 0.0$$026\pm 0.013$
Task prompt #2 Fixation-level $397\pm 0.012 0.0$$57\pm 0.009 0.0$$025\mathbf$$\pm 0.025$}
Word-level $045\pm 0.005 0.0$$43\pm 0.005 0.0$$026\pm 0.012$
Both fixation- and word-level $028\pm 0.003 0.0$$25\pm 0.004 0.0$$013\pm 0.009$
Question category (Kappa)
Task prompt #1 Fixation-level $\mathbf{0.299\pm 0.011}$ $074\pm 0.010 0.0$$045\pm 0.028$
Word-level $045\pm 0.009 0.0$$15\pm 0.009 0.0$$055\pm 0.023$
Both fixation- and word-level $035\pm 0.007 0.0$$06\pm 0.006 0.0$$014\pm 0.014$
Task prompt #2 Fixation-level $0.276\pm 0.011$ $\mathbf{0.107\pm 0.010}$ $\mathbf{0.067\pm 0.028}$
Word-level $052\pm 0.008 0.0$$37\pm 0.006 0.0$$031\pm 0.015$
Both fixation- and word-level $040\pm 0.007 0.0$$11\pm 0.006 0.0$$026\pm 0.015$
BLEU
Task prompt #1 Fixation-level $\mathbf{0.602\pm 0.006}$ $0.315\pm 0.003$ $0.314\pm 0.007$
Word-level $071\pm 0.001 0.0$$65\pm 0.001 0.0$$061\pm 0.002$
Both fixation- and word-level $069\pm 0.001 0.0$$63\pm 0.001 0.0$$058\pm 0.002$
Task prompt #2 Fixation-level $0.543\pm 0.006$ $\mathbf{0.332\pm 0.003}$ $\mathbf{0.336\pm 0.007}$
Word-level $054\pm 0.001 0.0$$48\pm 0.001 0.0$$047\pm 0.001$
Both fixation- and word-level $051\pm 0.001 0.0$$48\pm 0.000 0.0$$046\pm 0.001$
BERTScore
Task prompt #1 Fixation-level $\mathbf{0.774\pm 0.003}$ $0.631\pm 0.002$ $0.628\pm 0.004$
Word-level $0.418\pm 0.004$ $0.406\pm 0.003$ $0.394\pm 0.008$
Both fixation- and word-level $0.415\pm 0.004$ $0.403\pm 0.003$ $0.384\pm 0.007$
Task prompt #2 Fixation-level $0.744\pm 0.003$ $\mathbf{0.633\pm 0.001}$ $\mathbf{0.633\pm 0.004}$
Word-level $0.399\pm 0.003$ $0.385\pm 0.003$ $0.367\pm 0.008$
Both fixation- and word-level $0.400\pm 0.003$ $0.390\pm 0.002$ $0.383\pm 0.005$
QA accuracy
Task prompt #1 Fixation-level $\mathbf{76.3\pm 0.8}$ $\mathbf{65.1\pm 0.8}$ $65.2\pm 2.2$
Word-level $59.7\pm 0.8$ $56.9\pm 0.8$ $56.8\pm 2.0$
Both fixation- and word-level $61.8\pm 0.8$ $59.1\pm 0.8$ $57.3\pm 2.2$
[…]
Word-level $57.9\pm 0.8$ $6.2\pm 0.8 55$$5.1\pm 2.2$
Both fixation- and word-level $56.5\pm 0.7$ $3.9\pm 0.8 55$$5.2\pm 2.3$